OpenStudy (anonymous):

Given A (-6,4) and M(-1, -1) find B.

OpenStudy (anonymous):

what is B?

OpenStudy (anonymous):

thats all the information they gave me

OpenStudy (anonymous):

oh M is midpoint? if so then you know that if B has coordinates (x,y) you get \[\frac{-6+x}{2}=-1\] \[\frac{4+y}{2}=-1\]

OpenStudy (anonymous):

first equation give \[-x+6=-2\] \[x=8\]

OpenStudy (anonymous):

second equation gives \[4+y=-2\] \[y=-6\]

OpenStudy (anonymous):

so B is (8,-6)

OpenStudy (anonymous):

hold on i must have made a mistake somewhere

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

(-6, -4)

OpenStudy (anonymous):

first equation is \[-6+x=-2\] \[x=4\]

OpenStudy (anonymous):

B is (4,-6)